OpenStudy (anonymous):

My old nemesis... Any brave word-problem slayers know how to tackle this nightmare for me? Working together, two pumps can drain a certain pool in 4 hours. If it takes the older pump 9 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own?

OpenStudy (imstuck):

7.2 hours.

OpenStudy (anonymous):

Do you mind briefly explaining your logic?

OpenStudy (imstuck):

I do not mind at all!

OpenStudy (anonymous):

Gee thanks!

OpenStudy (imstuck):

This is a work problem. If the pumps do the job together in 4 hours, the ratio of hours to work is 1/4, meaning they can get 1/4 of the whole job done in one hour. If the old pump takes 9 hours to do the job alone, it can get 1/9 done in 1 hour. We need to find out how long it takes the new pump to do the job alone, so we will call that x.

OpenStudy (imstuck):

\[\frac{ 1 }{ 9 }+\frac{ 1 }{ x }=\frac{ 1 }{ 4 }\]

OpenStudy (imstuck):

to solve for x, you need to rid yourself of those pesky fractions, so we will multiply all the terms by the LCM of 36x. We then have this:

OpenStudy (imstuck):

4x + 36 = 9x

OpenStudy (anonymous):

Oh, thank you! I had set up that equation but subtracted 1/9 from 1/4 instead of doing the LCM.

OpenStudy (anonymous):

And got a funky answer..

OpenStudy (imstuck):

Can you see then how to solve for x?

OpenStudy (imstuck):

funky isn't good! ; )

OpenStudy (anonymous):

You are awesome on turning on light bulbs for people! Thank you for taking the time to enlighten me. Now I have actually learned something!

OpenStudy (imstuck):

lol

OpenStudy (imstuck):

Wow! Thank you so much for that!

OpenStudy (imstuck):

lol on the funky...

OpenStudy (anonymous):

Thanks for slaying my problem IMStuck :D

OpenStudy (imstuck):

You bet!

OpenStudy (imstuck):

lol again! Call me "The Slayer"...

OpenStudy (anonymous):

You can probably tell I like English way over math :/ But yes, its true. This math problem has been killed :D